# Explainable Deep Convolutional Multi-Type Anomaly Detection

### v1

@@ Abstract
Most explainable anomaly detection methods often identify anomalies but lack the capability to differentiate the type of anomaly. Furthermore, they often require the costly training and maintenance of separate models for each object category. The lack of specificity is a significant research gap, as identifying the type of anomaly (e.g., “Crack” vs. “Scratch”) is crucial for accurate diagnosis that facilitates cost-saving operational decisions across diverse application domains. While some recent large-scale Vision-Language Models (VLMs) have begun to address this, they are computationally intensive and memory-heavy, restricting their use in real-time or embedded systems. We propose MultiTypeFCDD, a simple and lightweight convolutional framework designed as a practical alternative for explainable multi-type anomaly detection. MultiTypeFCDD uses only image-level labels to learn and produce multi-channel heatmaps, where each channel is trained to correspond to a specific anomaly type. The model functions as a single, unified framework capable of differentiating anomaly types across multiple object categories, eliminating the need to train and manage separate models for each object category. We evaluated our proposed method on the Real-IAD dataset and it delivers results competitive with state-of-the-art complex models at significantly reduced parametric load and inference times. This makes it a highly practical and viable solution for real-world applications where computational resources are tightly constrained.
@@ 1 Introduction
Anomaly detection is a critical aspect in computer vision where the core objective is to identify patterns in visual data that do not conform to expected normal behaviour [pang_2021_deep_anomaly_detection_survey]. Recent years have witnessed significant adoption of deep learning-based anomaly detection in domains ranging from medical imaging [fernando2021deep], autonomous driving [bogdoll_2022_anomaly_driving_survey] to industrial manufacturing [alzarooni2025anomaly]. An important aspect of these sophisticated models is their explainability. Being able to understand why the model flagged something as an anomaly not only builds trust, but also helps the end users to make better and informed decisions, especially in sensitive areas like healthcare and autonomous driving [arrieta2020explainable, li2023survey]. By definition, anomalies are rare events, leading to a class imbalance where the normal samples vastly outnumber the abnormal ones [ruff_2021_anomaly_review]. Consequently, many state-of-the-art methods approach this problem of explainable anomaly detection from a one-class classification perspective [defard2021padim, deng2022anomaly, liu2023simplenet, liznerski_explainable_2021, roth2022towards].
A key limitation of all these methods is that they do not distinguish between different types of anomaly (unless separate models are trained for this purpose). This inhibits real-world decision making where understanding the type of anomaly is critical for determining the appropriate course of follow-up action [foorthuis2021nature]. For instance, in manufacturing metal parts, a system that can distinguish between a harmless discolouration and any rust or corrosion can trigger a simple cleaning process for the former, while initiating a more intensive grinding process or even scrapping the part entirely for the latter [bai2025comprehensive]. This characterisation is what enables an intelligent, cost-saving decision. Despite this clear need, the field of multi-type anomaly detection remains relatively unexplored [mokhtar2025detect].
@@
Recently, methods have begun to address this problem by exploiting vision-language models (VLMs) [mokhtar2025detect, sadikaj2025multiads] combining text prompts with visual features to segment multiple defect types. However, these models are computationally intensive, leading to substantial energy consumption and these translate into significant carbon emissions and financial costs [papa2024survey] Additionally, such architectures are often not well suited to systems with tight constraints on computational power and latency [marshall2025transformers], such as embedded systems, robotics and industrial automation. This highlights a clear need to explore alternative approaches that prioritise computational efficiency and practical deployability for real-world systems.
To address this research gap, our work proposes a novel deep convolutional multi-type anomaly detection method. We base the architecture on the Fully Convolutional Data Description (FCDD) [liznerski_explainable_2021] framework, extending it to produce multi-channel outputs, each corresponding to a distinct defect category, which we refer to as MultiTypeFCDD. We also develop a novel loss function to handle multiple types of anomaly. This enables the simultaneous detection and spatial localisation within an image for multiple pre-defined anomaly types. Our research addresses the general question: Given a pre-defined set of anomaly types found in various objects, can we uniquely identify these anomalies in any given object using a lightweight and efficient model? We demonstrate, on the publicly available Real-IAD dataset [wang2024real], that our proposed lightweight framework retains competitive performance to state of the art models while operating with significantly fewer parameters and lower inference times, making it a viable and practical solution for real-world applications with tight resource constraints.
@@ 2.2 Explainable Anomaly Detection and Segmentation
The black-box nature of deep learning models has driven the need to develop methods to aid transparency in the decision-making process [arrieta2020explainable]. Recent research have witnessed methods specifically designed for explainable anomaly detection, including feature-embedding methods such as PaDiM [defard2021padim], PatchCore [roth2022towards] and SoftPatch [jiang2022softpatch], and end-to-end trainable methods such as FCDD [liznerski_explainable_2021], SimpleNet [liu2023simplenet] and Reverse Distillation (RD) [deng2022anomaly].
@@ 2.3 Multi-Class vs Multi-Type Anomaly Detection
The methods discussed above frame anomaly detection from a one-class classification perspective where the goal is to separate out a single type of anomaly from normal data. This approach is sufficient for many anomaly detection benchmarks such as MVTec-AD [bergmann2021mvtec] and VisA [zou2022spot].
Recent literature has started to explore and use the terminology “multi-class anomaly detection” [he2024mambaad, jiang2024toward, you2022unified], which typically refers to identifying single types of anomaly across multiple objects. Additionally, most existing models are designed to segment anomalies within a single object, meaning that separate models might be required to process each object.
@@ 2.5 Research Gap and Our Contribution
The literature contains two main types of deep anomaly detectors: in one group, there are efficient, lightweight and inherently explainable anomaly detection models, but are typically limited to detecting only one type of anomaly (and multiple models are required to detect different types of anomaly). In the other group we have powerful VLMs which can perform multi-type anomaly detection and segmentation but are computationally expensive. This identifies a clear research gap: the need for a unified, lightweight and efficient model with inherent explainability capable of multi-type anomaly detection and segmentation.
We introduce a lightweight extension of the FCDD [28] framework called MultiTypeFCDD that generates multi-channel anomaly heatmaps, with each channel representing a specific defect category. This design allows for the simultaneous detection and localization of multiple predefined anomaly types.
@@ 3.2 Loss Function
The standard FCDD loss function is designed for binary anomaly detection where each image is either normal or anomalous. In our multi-type anomaly detection setting, each image may contain multiple anomaly classes. Therefore, we modify the FCDD loss function to accommodate each type of anomaly. The proposed objective function is defined as
@@ 4.1 Dataset
We follow the same train/test split procedure as in the original paper and use three training sets with noise ratios $\alpha\in\{0.1,0.2,0.4\}$, indicating the fraction of anomalous images in the training data.
@@ 4.2 Balanced Dataset Sampling: Redefining Epochs
Anomaly detection tasks inherently suffer from data imbalance, as datasets contain fewer anomalous samples than normal images. Given that we require images of various types of anomalies rather than “any” anomaly, the data imbalance problem becomes even more severe. Training a model on an imbalanced dataset can lead to overfitting to patterns in normal images and reduced sensitivity to anomalies [ghosh2024class, ruff_2021_anomaly_review]. To address this, we adopt a balanced sampling strategy during training, inspired by the sampling procedure for binary anomaly detection in [liznerski_explainable_2021] (Appendix E), where images from each anomaly class, including normal, are uniformly sampled per iteration.
While this improves performance and enhances the model’s ability to distinguish among various anomalies and normal images, it has an important computational implication: the number of iterations required to complete an epoch increases. Unlike a standard training epoch, which process each image exactly once, balancing requires the reshuffling and resampling of classes with fewer images, necessitating the analytical computation of the total number of iterations per epoch.
The problem is analogous to the generalised coupon collector problem [shank2013coupon], where each “coupon” corresponds to an image from a given class. Let each class $i\in\{1,2,\cdots,n\}$ have a target quota of $m_{i}$ (number of images per class) and is randomly chosen with equal probability. The expected number of iterations $E[T]$ required to see each class $i$ at least $m_{i}$ times is given by the formulation
$E[T]=n\int_{0}^{\infty}\left[1-\prod_{i=1}^{n}\big(1-S_{m_{i}}(t)e^{-t}\big)\right]dt,\quad$ (4)
where
$S_{m_{i}}(t)=\sum_{k=0}^{m_{i}-1}\frac{t^{k}}{k!}$
corresponds to the first $m_{i}$ terms of the series expansion of $e^{t}$.
For large quotas, direct computation of this integral is practically infeasible. Therefore in practice, one can use Monte Carlo simulation to estimate $E[T]$ empirically as
@@
where $T_{j}$ is the number of iterations in trial $j$ until all quotas are met, and $N_{trials}$ is the total number of Monte Carlo trials. This gives the empirical mean iterations per epoch, denoted as $\mu_{T}$. During training, images are processed in mini-batches. Therefore, the mean mini-batch iterations per epoch, which we refer to as balanced epochs, can be computed as $\mu^{batch}_{T}$.
Table 2 shows Monte Carlo simulation results (10000 iterations) on the three different Real-IAD training datasets across 9 image classes (8 anomaly classes and 1 normal class) at a mini-batch size of 32. When interpreted under the standard epoch definition (where each image is sampled exactly once), a balanced epoch corresponds to several standard epochs, e.g. at $\alpha=0.1$, one balanced epoch equals 8.41 standard epochs. Therefore, even when the model appears to converge in only a few balanced epochs, it has actually been exposed to the data multiple times, achieving better class coverage and more stable learning.
@@ 4.3.3 Hardware Configuration
4.4 Evaluation Metrics
We follow the Real-IAD [wang2024real] benchmark protocol, where model performance is evaluated at both the image and pixel levels. At the image level, we report metrics such as Image-level AUROC (I-AUROC), Average Precision (AP), and max F1-score (F1m) to assess performance across multiple defect types. At the pixel level, Pixel-level AUROC (P-AUROC) and Area Under the Per-Region Overlap (AUPRO) [bergmann2021mvtec] were computed. AUPRO captures region-wise localisation quality and penalises false detections on background areas, providing a robust measure of segmentation precision.
4.5 Benchmark Algorithms
We also report metrics for MultiTypeFCDD under a multi-class anomaly detection setting, because the modified framework with a single-channel anomaly output functions as a standard FCDD model that can be compared with state-of-the-art anomaly detectors, which identify the presence of anomalies but do not distinguish their types. Accordingly, we compare performance against PaDiM [defard2021padim], CFlow [gudovskiy2022cflow], PatchCore [roth2022towards], SimpleNet [liu2023simplenet], DeSTSeg [zhang2023destseg], RD [deng2022anomaly], and UniAD [you2022unified] and report I-AUROC and AUPRO metrics.
@@ 5.1 Multi-Type Anomaly Detection
The results of multi-type anomaly detection using our method are presented in Tables 3 and 4. At the image-level, mean I-AUROC reports over 94% across varying $\alpha$ values, followed by similar increasing trends for mean AP and F1m. This indicates the model’s capability to discriminate between normal and anomaly-specific images. At the pixel-level, the the model maintains robust localisation performance with mean P-AUROC above 92% and steady AUPRO improvements as $\alpha$ increases. In general, the model has strong capability of detecting various types of anomalies. However, it should be noted that certain anomaly types, such as Foreign Objects (YW) has lower performance values because the dataset contains fewer samples for this class compared to others and lacks consistent patterns when compared to classes like Scratch (HS). Figure 2 gives a visualisation of anomaly localisation across the eight defect classes for vaious objects at $\alpha=0.2$. We have also added a visual comparison of the anomaly heatmaps (see Fig.3) for various $\alpha$ values for a test object to illustrate the anomaly localisation performance of our model.
Figure 2: Illustration of multi-type anomaly detection across different objects ($\alpha=0.2$). Each pair shows the object (left) with the ground-truth anomaly and its corresponding anomaly heatmap (right) for each anomaly category, with colours for each anomaly type indicated in the legend.
Table 3: Image-Level Multi-Type Anomaly Detection Performance.
Anomaly Type $\alpha=0.1$ $\alpha=0.2$ $\alpha=0.4$
I-AUROC AP F1m I-AUROC AP F1m I-AUROC AP F1m
Pit 92.4 47.9 51.1 95.7 62.5 60.7 93.7 50.5 51.7
Deformation 99.4 79.7 72.4 99.7 90.7 84.6 99.6 88.5 81.7
Abrasion 97.4 71.4 69.7 98.7 74.0 69.0 99.0 77.7 74.1
Scratch 91.1 56.6 52.0 92.7 56.4 56.4 92.9 58.3 52.5
Damage 97.1 37.1 45.5 98.9 57.0 56.4 98.3 46.7 50.6
Missing Parts 89.9 56.9 53.9 91.0 63.5 59.2 90.7 53.0 51.9
Foreign Objects 98.2 21.0 37.3 99.0 37.7 46.3 99.2 57.3 52.9
Contamination 91.5 77.2 72.2 93.0 77.8 74.9 93.1 81.0 75.6
Mean 94.6 56.0 56.8 96.1 64.9 63.4 95.8 64.1 61.4
Table 4: Pixel-Level Multi-Type Anomaly Detection Performance.
Anomaly Type $\alpha=0.1$ $\alpha=0.2$ $\alpha=0.4$
P-AUROC AUPRO P-AUROC AUPRO P-AUROC AUPRO
Pit 92.7 72.0 94.5 77.0 95.1 74.8
Deformation 97.7 81.3 97.0 82.9 97.2 80.2
Abrasion 97.6 75.0 97.0 66.2 97.7 76.5
Scratch 84.6 64.2 86.5 65.1 86.8 65.7
Damage 92.0 53.3 93.1 70.5 94.9 59.5
Missing Parts 85.4 69.1 82.0 66.8 87.8 74.8
Foreign Objects 96.6 39.9 94.6 40.6 95.5 38.7
Contamination 90.3 75.7 91.2 78.1 92.1 80.0
Mean 92.1 66.3 92.0 68.4 93.4 68.8
Figure 3: Comparison of anomaly heatmaps for varying values of $\alpha$.

### v2

@@ Abstract
Explainable anomaly detection methods often have the capability to identify and spatially localise anomalies within an image but lack the capability to differentiate the type of anomaly. Furthermore, they often require the costly training and maintenance of separate models for each object category. The lack of specificity is a significant research gap because identifying the type of anomaly (e.g., “Crack” vs. “Scratch”) is crucial for accurate diagnosis that facilitates cost-saving operational decisions across diverse application domains. While some recent large-scale Vision-Language Models (VLMs) have begun to address this, they are computationally intensive and memory-heavy, restricting their use in real-time or embedded systems. We propose MultiTypeFCDD, a simple and lightweight convolutional framework designed as a practical alternative for explainable multi-type anomaly detection. MultiTypeFCDD uses only image-level labels to learn and produce multi-channel heatmaps, where each channel is trained to correspond to a specific anomaly type. The model functions as a single, unified framework capable of differentiating anomaly types across multiple object categories, eliminating the need to train and manage separate models for each object category. We evaluated our proposed method on the Real-IAD dataset and it delivers competitive results (96.4% I-AUROC) at just over 1% the size of state-of-the-art VLM models used for similar tasks. This makes it a highly practical and viable solution for real-world applications where computational resources are tightly constrained.
@@ 1 Introduction
Anomaly detection is a critical aspect in computer vision where the core objective is to identify patterns in visual data that do not conform to expected normal behaviour [pang_2021_deep_anomaly_detection_survey]. Recent years have witnessed significant adoption of deep learning-based anomaly detection in domains ranging from medical imaging [fernando2021deep], autonomous driving [bogdoll_2022_anomaly_driving_survey] to industrial manufacturing [alzarooni2025anomaly]. An important aspect of these sophisticated models is their explainability – by explainability in this paper, we mean spatially localising the anomaly within the input image. Being able to understand why the model flagged something as an anomaly not only builds trust, but also helps the end users to make better and informed decisions, especially in sensitive areas like healthcare and autonomous driving [arrieta2020explainable, li2023survey]. By definition, anomalies are rare events, leading to a class imbalance where the normal samples vastly outnumber the abnormal ones [ruff_2021_anomaly_review]. Consequently, many state-of-the-art methods approach this problem of explainable anomaly detection from a one-class classification perspective [defard2021padim, deng2022anomaly, liu2023simplenet, liznerski_explainable_2021, roth2022towards].
A key limitation of existing anomaly detection methods is that they often require training separate models for each object class, which limits scalability and practicality in real-world applications, as illustrated in Fig. 1(a). Even “multi-class anomaly detection” approaches address this only partially [you2022unified], as unified models can detect anomalies but do not distinguish between different types of anomaly, as shown in Fig. 1(b). This inhibits real-world decision making where understanding the type of anomaly is critical for determining the appropriate course of follow-up action [foorthuis2021nature].
@@
Recently, methods have begun to address this problem by exploiting vision-language models (VLMs) [mokhtar2025detect, sadikaj2025multiads], combining text prompts with visual features to segment multiple defect types. However, these models are computationally intensive, leading to substantial energy consumption and which translates into significant carbon emissions and financial costs [papa2024survey] Additionally, such architectures are often not well-suited to systems with tight constraints on computational power and latency [marshall2025transformers], such as embedded systems, robotics and industrial automation. This highlights a clear need to explore alternative approaches that prioritise computational efficiency and practical deployability.
To address this research gap, our work proposes a novel deep convolutional multi-type anomaly detection method. We base the approach on the Fully Convolutional Data Description (FCDD) [liznerski_explainable_2021]. This enables the simultaneous detection and spatial localisation within an image for multiple pre-defined anomaly types. Our research makes three main contributions:
  • We contribute a novel fully convolutional model for multi-type anomaly detection with multiple output channels sensitised to different types of anomaly, which we refer to as MultiTypeFCDD.
  • We contribute a novel loss function and training framework for multi-type anomaly detection, extending the original FCDD framework.
  • We demonstrate, on the publicly available Real-IAD dataset [wang2024real], that our proposed lightweight framework retains competitive performance to state-of-the-art models while operating with significantly fewer parameters and lower inference times.
The paper is structured as follows. In Section 2, we review related work to explainable anomaly detection. In Section 3, we describe the proposed method for multi-type anomaly detection. In Section 4, we describe the experimental setup to evaluate the proposed method, including the publicly available RealIAD dataset and benchmark algorithms. In Section 5, we provide the results and in Section 6, we conclude the work.
Figure 1: Illustration of multi-type anomaly detection on two objects (USB and Woodstick) from the Real-IAD dataset [wang2024real]. (a) In multi-class anomaly detection, most explainable methods require training on separate networks to detect anomalies in each object class. (b) Recent research has introduced unified frameworks to detect anomalies across multiple objects. However, both (a) and (b) are only capable of detecting an anomaly and not what type of anomaly. (c) Our proposed framework enables the detection of multiple types of anomalies occurring across diverse objects using a single, unified model.
@@ 2.2 Explainable Anomaly Detection and Segmentation
The black-box nature of deep learning models has driven the need to develop methods to aid transparency in the decision-making process [arrieta2020explainable]. Several techniques have been specifically designed for explainable anomaly detection, including feature-embedding methods such as PaDiM [defard2021padim], PatchCore [roth2022towards] and SoftPatch [jiang2022softpatch], and end-to-end trainable methods such as FCDD [liznerski_explainable_2021], SimpleNet [liu2023simplenet] and Reverse Distillation (RD) [deng2022anomaly].
@@ 2.3 Multi-Class vs Multi-Type Anomaly Detection
The methods discussed above frame anomaly detection from a one-class classification perspective, where the goal is to learn the representation of the normal data, and thereby separate everything anomalous as a single, separate class. This approach is sufficient for many anomaly detection benchmarks such as MVTec-AD [bergmann2021mvtec] and VisA [zou2022spot].
Recent literature has started to explore and use the terminology “multi-class anomaly detection” [he2024mambaad, you2022unified], which typically refers to identifying a single anomaly class across multiple objects. Additionally, most existing models are designed to segment anomalies within a single object, meaning that separate models might be required to process each object.
@@ 2.5 Research Gap and Our Contribution
The literature contains two main types of deep anomaly detectors: in one group, there are lightweight anomaly detection models, based on convolutional networks, which are typically limited to detecting only one type of anomaly. In the other group, we have powerful VLMs which can perform multi-type anomaly detection and segmentation but are computationally expensive. This identifies a clear research gap: the need for a unified, lightweight model with inherent explainability, capable of multi-type anomaly detection and segmentation.
We introduce an extension of the FCDD [28] framework, which we term MultiTypeFCDD, that generates multi-channel anomaly heatmaps, with each channel representing a specific type of anomaly. This design enables the simultaneous detection and localization of multiple predefined anomaly types.
@@ 3.2 Loss Function
The standard FCDD loss function is designed for binary anomaly detection where each image is either normal or anomalous. In our multi-type anomaly detection setting, we extend the FCDD loss function to accommodate each type of anomaly. The model must be trained under two conditions: 1. there must examples of anomalies available (so it cannot be trained as a one-class classifier unlike the original FCDD) and 2. the training data set must contain images with single anomaly types (although at inference time the model can predict multiple types of anomaly in a single image).
The proposed objective function is defined as
@@ 4.1 Dataset
We follow the same train/test split procedure as in the original paper [wang2024real] and use three training sets with noise ratios
$\alpha\in\{0.1,0.2,0.4\}$ (4)
which is a parameter that controls the number of anomalous samples mixed into the training data.
@@ 4.2 Balanced Dataset Sampling: Redefining Epochs
Anomaly detection tasks inherently suffer from data imbalance, as datasets contain fewer anomalous samples than normal images. Given that we require images of various types of anomalies rather than “any” anomaly, the data imbalance problem becomes even more severe. Training a model on an imbalanced dataset can lead to overfitting to patterns in normal images and reduced sensitivity to anomalies [ruff_2021_anomaly_review]. To address this, we adopt a balanced sampling strategy during training, inspired by the sampling procedure for binary anomaly detection in [liznerski_explainable_2021], where images from normal and each anomaly class are uniformly sampled per iteration.
While this improves performance and helps the model better distinguish between normal and various anomaly types, it has an important computational implication: the number of iterations required to complete an epoch increases. Unlike standard training epochs that process each image exactly once, balancing requires the reshuffling and resampling under-represented classes, necessitating the analytical computation of the total number of iterations per epoch.
The problem is analogous to the generalised coupon collector problem [shank2013coupon], where each “coupon” corresponds to an image from a given target class. Let each class $i\in\{1,2,\cdots,n\}$ have a target quota of $m_{i}$ (number of images per class) and is randomly chosen with equal probability. During sampling, if any class $i$ has reached its full quota $m_{i}$, its images are reshuffled and can be sampled again in subsequent iterations. Let $E[T]$ be the expected number of iterations needed to sample at least $m_{i}$ images from every class. Although a closed-form expression for $E[T]$ is given in [shank2013coupon], direct computation is practically infeasible, especially when $m_{i}$ is large $(m_{i}\gg 1)$ as it involves computing factorials of $m_{i}$. Therefore, in practice, one can use Monte Carlo simulation to estimate $E[T]$ empirically as
@@
where $T_{j}$ is the number of iterations in trial $j$, and $N_{trials}$ is the total number of Monte Carlo trials. This gives the empirical mean iterations per epoch, $\mu_{T}$. During training, images are processed in mini-batches. Therefore, the mean number of mini-batch iterations per epoch, which we refer to as balanced epochs, can be computed as $\mu^{batch}_{T}$.
Table 2 shows Monte Carlo simulation results (10000 trials) on the three different Real-IAD training datasets across 9 image classes (8 anomaly classes and 1 normal class) at a mini-batch size of 32. When interpreted under the standard epoch definition (where each image is sampled exactly once), a balanced epoch corresponds to several standard epochs, e.g. at $\alpha=0.1$, one balanced epoch equals 8.41 standard epochs. Therefore, even when the model appears to converge in only a few balanced epochs, it has actually seen the data multiple times, achieving better class coverage and more stable learning.
@@ 4.3.3 Hardware Configuration:
4.4 Metrics
We follow the Real-IAD [wang2024real] benchmark protocol, where model performance is evaluated at both the image and pixel levels in terms of the Receiver Operating Characteristic (ROC) curve, defined by the True Positive Rate (TPR) versus the False Positive Rate (FPR). We summarise the ROC curve using the area under the ROC curve (AUROC) for both the Image-level AUROC (I-AUROC) and Pixel-level AUROC (P-AUROC). For image-level metrics, we compute the anomaly score as the mean value of the anomaly heatmap per channel as given in Eq. 2.
To assess the performance of spatial localisation of the anomaly, we calculate the per-region overlap (PRO) score [bergmann2021mvtec]. Let $C_{i,k}$ represent the set of pixels for the $k^{th}$ ground-truth component in image $i$, and $P_{i}(\tau)$ be the set of predicted anomalous pixels for image $i$ at a given threshold $\tau$; the PRO score is defined as the average coverage of these components across the entire dataset,
$\text{PRO}(\tau)=\frac{1}{N}\sum_{i}\sum_{k}\frac{|P_{i}(\tau)\cap C_{i,k}|}{|C_{i,k}|}$ (6)
where $N$ denotes the total count of ground-truth components in the dataset. We use the Area Under the PRO curve (AUPRO) to summarise the PRO score in a single metric, which is defined as the normalised area under the PRO-FPR curve up to an FPR limit of 0.3.
4.5 Benchmarking
To benchmark MultiTypeFCDD, we compare it to several other methods that perform explainable anomaly detection (using the Real-IAD dataset), including PaDiM [defard2021padim], CFlow [gudovskiy2022cflow], PatchCore [roth2022towards], SimpleNet [liu2023simplenet], DeSTSeg [zhang2023destseg], RD [deng2022anomaly], and UniAD [you2022unified]. These benchmark methods do not distinguish types of anomaly, so we only compared performance by treating all anomaly types as a single class. Although MultiTypeFCDD uses weak image-level labels during training, these do not provide spatial supervision, so comparison with unsupervised one-class methods remains fair. To handle multiple objects, we compared multiple models trained on separate objects (where necessary). We compared performance in terms of I-AUROC to characterise anomaly detection at the image level, and AUPRO to characterise the success of spatial localisation of the anomaly.
@@ 5.1 Multi-Type Anomaly Detection
The results of multi-type anomaly detection using our method are presented in Table 3. At the image-level, mean I-AUROC reports over 95% across varying $\alpha$ values, indicating the model’s capability to discriminate between normal and anomaly-specific images. At the pixel level, the model maintains robust localisation performance with a mean P-AUROC above 93% and steady AUPRO improvements as $\alpha$ the number of iterations increases. In general, the model has a strong capability of detecting various types of anomalies. However, certain anomaly types, such as Foreign Objects (YW), exhibit lower performance because they have fewer samples, fewer ground-truth pixels and less consistent patterns due to their high semantic diversity when compared to classes like Deformation (BX).
Figure 3 gives a visualisation of anomaly localisation across the eight defect classes for various objects at $\alpha=0.2$. Because the Real-IAD dataset contains only one defect type per image, Figure 4 includes custom-generated examples created by merging two defect instances of the same object class into a single sample. This example demonstrates that the model is capable of identifying and localising multiple anomalies within one image. Furthermore, to illustrate the localisation behaviour across multiple $\alpha$ values, Fig. 5 provides a visual comparison of anomaly heatmaps for a representative test object.
Figure 3: Examples of multi-type anomaly detection across multiple objects ($\alpha=0.2$). Each pair of images in the grid shows the input image with the ground-truth anomaly highlighted by the circle (left) and the corresponding predicted anomaly heatmap by MultiTypeFCDD overlaid on the input image (right), with colours defining each anomaly type indicated in the legend on the right.
Figure 4: Example of multi-type anomaly detection using manually edited synthetic test images containing multiple co-occurring anomalies ($\alpha=0.2$). For each image, the corresponding overlaid anomaly heatmap highlights all detected anomalous regions.
Table 3: Image and pixel-level multi-type anomaly detection performance across different $\alpha$ levels: 0.1, 0.2, 0.4.
Anomaly Type I-AUROC P-AUROC AUPRO
0.1 0.2 0.4 0.1 0.2 0.4 0.1 0.2 0.4
Pit (AK) 95.2 95.7 96.2 97.2 94.8 97.0 84.8 80.7 85.6
Deformation (BX) 99.6 99.3 99.4 98.4 96.3 96.1 87.4 86.4 87.4
Abrasion (CH) 98.3 98.2 99.1 96.0 95.5 96.4 81.8 84.0 83.7
Scratch (HS) 90.0 94.4 95.9 85.3 90.5 88.9 64.0 72.5 67.6
Damage (PS) 98.7 98.6 98.8 96.0 96.4 97.8 79.2 88.3 88.7
Missing Parts (QS) 92.0 94.3 96.1 84.2 91.5 91.8 70.6 82.3 80.5
Foreign Objects (YW) 98.7 99.3 99.4 95.0 93.3 91.5 57.5 60.3 49.2
Contamination (ZW) 92.9 93.2 95.0 95.6 95.8 96.6 79.3 81.9 82.3
Mean 95.7 96.6 97.5 93.5 94.3 94.5 75.6 79.6 78.1
Figure 5: Comparison of anomaly heatmaps for a test object (audiojack) across different $\alpha$ levels: 0.1, 0.2, 0.4.